# CountZES: Counting via Zero-Shot Exemplar Selection

### v1

@@ Abstract
Object counting in complex scenes remains challenging, particularly in the zero-shot setting, where the goal is to count instances of unseen categories specified only by a class name. Existing zero-shot object counting (ZOC) methods that infer exemplars from text either rely on open-vocabulary detectors, which often yield multi-instance candidates, or on random patch sampling, which fails to accurately delineate object instances. To address this, we propose CountZES, a training-free framework for object counting via zero-shot exemplar selection. CountZES progressively discovers diverse exemplars through three synergistic stages: Detection-Anchored Exemplar (DAE), Density-Guided Exemplar (DGE), and Feature-Consensus Exemplar (FCE). DAE refines open-vocabulary detections to isolate precise single-instance exemplars. DGE introduces a density-driven, self-supervised paradigm to identify statistically consistent and semantically compact exemplars, while FCE reinforces visual coherence through feature-space clustering. Together, these stages yield a diverse, complementary exemplar set that balances textual grounding, count consistency, and feature representativeness. Experiments on diverse datasets demonstrate CountZES superior performance among ZOC methods while generalizing effectively across natural, aerial and medical domains. Code is available at GitHub.
@@ 1 Introduction
Figure 1: Zero-shot object counting comparison of our CountZES against recently proposed T2ICount [qian2025t2icount] and GeCo [pelhan2024novel] across diverse benchmarks spanning natural, aerial, and medical domains.
@@
Object counting plays a vital role in a wide range of applications, including crowd and traffic analysis [khan2023revisiting], agriculture [farjon2023deep], and biomedical microscopy [singh2024systematic]. Traditional counting approaches are typically class-specific, tailored for predefined categories such as crowds [babu2022completely], vehicles [li2020bi], or cells [tyagi2023degpr]. However, these methods require retraining whenever a new class is introduced, limiting their scalability. To overcome this, class-agnostic counting has emerged, aiming to generalize to unseen categories [hui2024class, yang2021class, huang2023interactive]. Existing class-agnostic methods can be categorized into three paradigms: few-shot object counting (FSOC), reference-less counting (RLC), and zero-shot object counting (ZOC).
@@
FSOC methods rely on a small set of annotated exemplars provided at inference [yang2021class, nguyen2022few, he2024few], which limits scalability [zhai2024zero, zhang2025enhancing]. In contrast, RLC estimates counts directly from global visual cues [hobley2022learning, liu2022countr, d2024afreeca, ranjan2021learning], but lacks semantic awareness and tends to focus on visually dominant patterns [zhu2024zero, xu2023zero]. To address these issues, ZOC has emerged as a promising direction [xu2023zero, zhu2024zero, kang2024vlcounter, jiang2023clip], enabling the counting of arbitrary unseen categories from text prompts alone.
@@
Current ZOC methods fall into three main streams. (1) Image–text interaction–based methods leverage vision–language models (VLMs) to align text and image embeddings for density regression. While scalable to unseen classes, they often fail to capture fine-grained visual details, especially for atypical objects [jiang2023clip]. (2) Class-relevant exemplar selection discover visual patches that align with the target class by comparing arbitrarily sampled regions with class prototypes [xu2023zero], but random patch sampling often produces partial or noisy exemplars. (3) To overcome this, detection-driven exemplar selection employs open-vocabulary detectors (OVD) to propose text-conditioned boxes as candidate exemplars. However, OVDs often fail to provide clean single-instance exemplars, especially in dense scenes. For example, VA-Count [zhu2024zero] employs GroundingDINO [liu2024grounding] to extract candidate exemplars and then uses a binary filter to verify single-object instances. However, this pipeline struggles when GroundingDINO provides few or no single-instance boxes. As shown in Fig. 2(a), for the class ’apple’, only a handful of valid single-instance boxes are detected, leading to an extremely limited exemplar pool. Conversely, for the category ’legos’, it predicts only one large box encompassing all instances. This dependence on the detector’s initial proposals creates a single point of failure that propagates through the counting pipeline. Furthermore, Fig. 2(b) illustrates that counting accuracy is highly sensitive to exemplar choice, rendering exemplar selection as a critical bottleneck. These limitations underscore the need for a principled exemplar selection strategy for ZOC.
To address this need, we introduce CountZES, a training-free framework for object counting via zero-shot exemplar selection, Fig. 2(c). CountZES formulates exemplar discovery as a progressive, multi-stage process comprising Detection-Anchored Exemplar (DAE), Density-Guided Exemplar (DGE), and Feature-Consensus Exemplar (FCE) stages. Each stage contributes an exemplar, resulting in a complementary set that balances textual alignment, count-driven consistency, and feature-level representativeness, respectively. The first stage, DAE, begins with a text-conditioned CLIP similarity map [radford2021learning] and GroundingDINO detections. To overcome the detector’s limited ability to provide clean single-instance boxes, it employs the proposed Similarity-guided SAM-based Exemplar Selection (SSES), which refines coarse multi-instance detections using CLIP-guided point prompts and entropy-based optimization, isolating localized, text-aligned single-object regions. The second stage, DGE, shifts from text grounding to a density-driven self-supervised paradigm. It generates a DAE-conditioned density map and leverages Peak-to-Point (P2P) prompting for candidate region discovery, followed by RoI-based single-instance filtering to retain boxes whose integrated density corresponds to a single object. The most reliable exemplar is then selected through the Pseudo-GT Guided Exemplar Selection (GGES) module, which balances count consistency, relative to statistically inferred pseudo-ground truth (GT) count, with semantic compactness. Finally, FCE emphasizes feature-level coherence. Using the DGE exemplar as a reference, the Feature-based Representative Exemplar Selection (FRES) module projects each candidate onto SAM’s feature space and identifies most representative exemplar via clustering.
Figure 2: (a) Detectors like GroundingDINO [liu2024grounding] often yield very few or even fail to provide single instance boxes. (b) Counting is sensitive to exemplar choice. (c) Introducing CountZES, a training-free, multi-stage exemplar selection approach for ZOC.
@@
Importantly, CountZES is a training-free framework that operates without any task-specific fine-tuning. While it leverages pretrained VLM, segmentation, and counting models, its exemplar selection pipeline remains independent of any dataset-specific optimization. This design enables improved generalization across datasets which is validated through extensive evaluations across diverse benchmarks spanning natural, aerial and medical domains, Fig. 1. In summary, we make following contributions:
@@
  • We propose CountZES, a training-free framework for object counting via zero-shot exemplar selection across three synergistic stages (DAE, DGE, and FCE) that collectively balance textual grounding, count consistency, and feature representativeness, yielding a diverse set.
@@
  • CountZES introduces dedicated modules at each stage, with SSES refining multi-instance detections into precise single-instance boxes. P2P prompting and RoI count-based filtering for density-driven exemplar discovery, and GGES and FRES for exemplar refinement through pseudo GT guidance and feature-level consensus, respectively.
@@
  • Being training-free, CountZES demonstrates better generalization across datasets and even across domains. Extensive experiments across natural, aerial, and medical cell-counting benchmarks demonstrate its robustness and superior performance compared to SOTA methods.
@@ 2 Related Work
To eliminate the dependence on manually annotated exemplars in FSOC methods [lu2018class, ranjan2021learning, liu2022countr, djukic2023low, he2024few], Xu et al. [xu2023zero] introduce the ZOC task, which estimates object counts from a class name by employing a text-conditioned variational autoencoder (VAE) that synthesizes visual exemplars through patch–prototype pseudo-labels. However, the arbitrary selection of patches hinders precise object delineation, thereby degrading exemplar quality. Subsequent works [jiang2023clip, kang2024vlcounter, amini2023open] leverage image–text alignment to model object–class correlations without relying on visual exemplars, enhancing scalability but often failing to preserve structural detail for irregularly shaped objects. Detection-driven approaches such as DAVE [pelhan2024dave] adopt a detect-and-verify pipeline but bias toward majority-class counts. Recent VLM-based frameworks like PseCo [huang2024point] employs a three-stage point, segment and count design integrating SAM for segmentation and CLIP for verification, while GeCo [pelhan2024novel] unifies detection, segmentation, and counting with a single-stage dense query encoder–decoder architecture. VA-Count [zhu2024zero] instead leverages GroundingDINO for text-guided detection and performs a simple logits-based exemplar selection (top-3), yielding limited exemplar diversity and relying on DINO’s initial proposals (often multi-instance boxes). Our CountZES explicitly handles the challenge of limited or no single instance boxes coming from the detector and progressive multi-stage exemplar selection offers a natural diversity in the exemplar set. Finally, T2ICount [qian2025t2icount] exploits text-to-image diffusion priors for open-world counting but may exhibit dataset bias due to training. In contrast, our CountZES being training-free demonstrates superior generalization ability.
@@ 3 Method
Figure 3: Overview of CountZES. The pipeline operates in three stages. DAE refines GroundingDINO detections using CLIP similarity via SSES to obtain text-aligned single-instance regions. DGE then adopts a density-driven paradigm: it generates a density map conditioned on DAE exemplar, uses P2P prompting to identify candidate exemplars, applies RoI-based single-instance filtering, and selects the most reliable exemplar through GGES. Finally, FCE focuses on visual coherence by projecting candidate exemplars onto SAM’s feature map and using FRES to select the representative exemplar. An exemplar from each stage form a diverse exemplar set for final object count.
We introduce CountZES, a training-free framework for ZOC (Fig. 3), which progressively discovers high-quality exemplars through three synergistic stages, namely DAE (Sec. 3.1), DGE (Sec. 3.2), and FCE (Sec. 3.3). Each stage contributes a distinct, yet complementary, perspective to exemplar selection as discussed in subsequent sections.
@@ 3.1 Detection Anchored Exemplar (DAE)
Initial Box Selection: Each detection box $\mathbf{b}_{i}$ serves as a candidate region, within which $S$ provides fine-grained confidence cues about semantic alignment with $t$. However, not all detections correspond to clean, single-instance regions, as some may encompass multiple instances or include background areas. To quantify the internal consistency of each region, we require a measure that captures how concentrated or dispersed the similarity responses are within the box. We therefore compute the entropy of the CLIP similarity distribution inside $\mathbf{b}_{i}$, which reflects the spatial uniformity of similarity values. Intuitively, boxes that include multiple objects or background exhibit high entropy, as similarity fluctuates across different semantic regions and object boundaries. In contrast, a box that tightly encloses a single, well-aligned instance yields low entropy, indicating compact and coherent similarity responses in $S$.
@@
where $\alpha=0.5$ controls the trade-off between semantic confidence and spatial purity, the latter reflecting how consistent $S$ is within $\mathbf{b}_{i}$. Higher $J_{\text{box}}(\mathbf{b}_{i})$ indicates boxes that are both confident (high $c_{i}$) and internally consistent (low $\widehat{H}$), favoring single-instance regions over cluttered ones. The top-scoring box, $\mathbf{b}^{c}=\arg\max_{\mathbf{b}_{i}}J_{\text{box}}(\mathbf{b}_{i})$, is selected as the coarse region of interest. Why coarse? Although this entropy-weighted scoring favors regions with fewer overlapping instances or backgrounds, it does not guaranty that the selected box strictly contains a single instance. When all GroundingDINO detections correspond to grouped regions, even the top-scoring box may include multiple instances. Hence, this coarse selection is refined next by the similarity-guided SAM-based exemplar selection (SSES) module to obtain an instance-specific region.
Sim-guided SAM-based Exemplar Selection (SSES): Given the coarse region $\mathbf{b}^{c}$, SSES refines it into a clean, instance-specific exemplar by leveraging the similarity map $S$ and SAM’s segmentation capability. The core intuition is that regions containing well-aligned object instances will exhibit distinct peaks in $S$, representing local maxima of semantic alignment with the target text. To identify high-similarity regions, we sample $k$ peaks from $S$ within $\mathbf{b}^{c}$ by applying percentile-based threshold rather than a fixed similarity cutoff. Percentiles are computed over the full similarity map $S$, normalizing thresholds to each image’s global similarity statistics. Starting from a high percentile (e.g., 90th), the threshold is gradually relaxed in fixed steps ($\Delta p=10$) until at most $k=16$ peaks (see supplementary) are detected within $\mathbf{b}^{c}$. This adaptive relaxation dynamically adjusts the confidence level for peak selection based on the strength and density of responses in the coarse region. Formally, for a percentile $p$, the threshold is defined as $\tau(p)=\operatorname{Perc}_{p}(S)$, and we retain spatial locations inside $\Omega(\mathbf{b}^{c})=\{(x,y)\,|\,x_{0}\leq x\leq x_{1},\,y_{0}\leq y\leq y_{1}\}$ whose similarity values exceed this threshold:
@@
where $w_{\text{sim}}=w_{\text{ent}}=0.5$. The optimal mask is selected as $M_{\text{DAE}}=\arg\max_{M_{i}}J_{\text{mask}}(M_{i})$, and its corresponding bounding box ($\mathbf{b}_{\text{DAE}}$) defines the refined, instance-specific region. In essence, SSES acts as a fine-grained refinement stage, transforming the coarse detector-derived region $\mathbf{b}^{c}$ into a single-instance exemplar $\mathbf{b}_{\text{DAE}}$. When the coarse region already corresponds to a confident single instance detection, the sampled point prompts within it consistently produce the same mask, leading SSES to retain the original region. In such cases, SSES effectively verifies rather than alters the Grounding DINO detection, confirming it as a reliable exemplar. The resulting DAE thus provides a semantically grounded and spatially refined support region used in subsequent stages of exemplar selection and counting.

### v2

@@ Abstract
Object counting in complex scenes is particularly challenging in the zero-shot (ZS) setting, where instances of unseen categories are counted using only a class name. Existing ZS counting methods that infer exemplars from text often rely on off-the-shelf open-vocabulary detectors (OVDs), which in dense scenes suffer from semantic noise, appearance variability, and frequent multi-instance proposals. Alternatively, random image-patch sampling is employed, which fails to accurately delineate object instances. To address these issues, we propose CountZES, an inference-only approach for object counting via ZS exemplar selection. CountZES discovers diverse exemplars through three synergistic stages: Detection-Anchored Exemplar (DAE), Density-Guided Exemplar (DGE), and Feature-Consensus Exemplar (FCE). DAE refines OVD detections to isolate precise single-instance exemplars. DGE introduces a density-driven, self-supervised paradigm to identify statistically consistent and semantically compact exemplars, while FCE reinforces visual coherence through feature-space clustering. Together, these stages yield a complementary exemplar set that balances textual grounding, count consistency, and feature representativeness. Experiments on diverse datasets demonstrate CountZES superior performance among ZOC methods while generalizing effectively across domains. Code is available at GitHub.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Zero-shot object counting comparison of our CountZES against T2ICount (Qian et al., 2025) , GeCo (Pelhan et al., 2024a), and TFOC (Shi et al., 2024) across diverse benchmarks spanning natural, aerial, and medical domains.
@@
Object counting plays a vital role in a wide range of applications. Traditional counting approaches are typically class-specific, tailored for predefined categories such as crowds (Babu Sam et al., 2022), vehicles (Li et al., 2020), or cells (Tyagi et al., 2023). However, these methods require retraining whenever a new class is introduced, limiting their scalability. To overcome this, class-agnostic counting has emerged, aiming to generalize to unseen categories (Yang et al., 2021; Huang et al., 2023; Hui et al., 2024). Existing class-agnostic methods can be categorized into three paradigms: few-shot object counting (FSOC), reference-less counting (RLC), and zero-shot object counting (ZOC).
@@
FSOC methods rely on a small set of annotated exemplars provided at inference, which limits scalability (Zhai et al., 2024; Zhang et al., 2025). In contrast, RLC estimates counts directly from global visual cues (Hobley and Prisacariu, 2023; D’Alessandro et al., 2024), but lacks semantic awareness and tends to focus on visually dominant patterns (Xu et al., 2023; Zhu et al., 2024). To address these issues, ZOC has emerged as a promising direction (Jiang et al., 2023; Xu et al., 2023; Kang et al., 2024), enabling the counting of arbitrary unseen categories from text prompts alone.
@@
Current ZOC methods fall into three main streams. (1) Image–text interaction–based methods leverage vision–language models (VLMs) to align text and image embeddings for density regression. While scalable to unseen classes, they often fail to capture fine-grained visual details, especially for atypical objects (Jiang et al., 2023). (2) Class-relevant exemplar selection discover visual patches that align with the target class by comparing arbitrarily sampled regions with class prototypes (Xu et al., 2023), but random sampling often produces partial or noisy exemplars. (3) To overcome this, detection-driven exemplar selection employs open-vocabulary detectors (OVD) to propose text-conditioned boxes as candidate exemplars. However, OVDs often fail to provide clean single-instance exemplars, especially in dense scenes. For example, VA-Count (Zhu et al., 2024) employs GroundingDINO (Liu et al., 2024) to extract candidate exemplars and then uses a binary filter to verify single-object instances. However, this pipeline struggles when GroundingDINO provides few or no single-instance boxes, Figure. 2(a). Furthermore, counting accuracy is sensitive to exemplar choice and randomly selected exemplars are often insufficiently representative of the target instances, leading to underestimated object counts, Figure. 2(b).
Despite its importance, exemplar selection in ZOC remains largely underexplored. Existing methods, such as ZSOC (Xu et al., 2023) and VA-Count, rely on random patch sampling or simple confidence-based filtering of OVD detections, respectively. However, object counting is inherently a dense prediction task that poses challenges fundamentally different from those in sparse settings. Heavy occlusion, scale variation, and tightly packed instances make patch-level sampling or off-the-shelf detection unreliable, Figure. 2(a). A natural alternative would be to train dedicated dense detectors like GeCo (Pelhan et al., 2024a). However, such approaches require costly box-level annotations in dense scenes and often internalize dataset-specific priors, resulting in limited cross-domain generalization (see GeCo in Figure. 1). Moreover, retraining or finetuning off-the-shelf vision backbones for each new domain not only requires large-scale annotated training data but is computationally expensive and impractical in real-world settings. This motivates the need for approaches that can fully exploit the rich representations of pretrained foundation models like CLIP (Radford et al., 2021) and SAM (Kirillov et al., 2023) while avoiding task-specific supervision. Existing inference-only approaches, such as TFOC (Shi et al., 2024) and TFCAC (Lin et al., 2025), primarily treat the ZOC task as a segmentation problem, relying solely on SAM, which exhibits limited performance in dense and congested scenes (Ma et al., 2023). Although exemplar-based density estimation has long been recognized as highly effective for dense counting scenarios (Liu et al., 2018), to our knowledge, no existing inference-only approach targets exemplar selection in ZOC.
@@
To fill this gap, we introduce CountZES, an inference-only pipeline for object counting via zero-shot exemplar selection, Figure. 2(c). CountZES formulates exemplar discovery as a multi-stage process comprising Detection-Anchored Exemplar (DAE), Density-Guided Exemplar (DGE), and Feature-Consensus Exemplar (FCE) stages. Each stage contributes an exemplar, resulting in a complementary set that balances textual alignment, count-driven consistency, and feature-level representativeness, respectively. We thus cast zero-shot counting as an exemplar inference problem under semantic, statistical, and feature-level priors. Importantly, CountZES adopts an inference-only exemplar selection pipeline that operates without any ZOC-specific fine-tuning. While the framework leverages pretrained VLM, segmentation, and density-estimation components for semantic alignment and counting, these models remain fixed and are not optimized within CountZES. This design enables improved generalization, which is validated through extensive evaluations across diverse benchmarks spanning natural, aerial, and medical domains (Figure. 1). In summary, we make following core contributions:
@@
  • We propose CountZES, an inference-only pipeline for ZOC via principled exemplar selection across three stages (DAE, DGE, and FCE) that collectively balance textual grounding, count consistency, and feature representativeness, yielding a diverse exemplar set.
@@
  • CountZES introduces dedicated modules at each stage: Similarity-guided SAM-based Exemplar Selection (SSES) refines coarse multi-instance OVD detections into precise single-instance exemplars; Peak-to-Point (P2P) prompting with RoI count-based filtering enables density-driven exemplar discovery; and Pseudo-GT Guided Exemplar Selection (GGES) together with Feature-based Representative Exemplar Selection (FRES) refines exemplars via statistical consistency and feature-level consensus.
@@
  • Being an inference-only exemplar selection approach, CountZES avoids ZOC-specific optimization and demonstrates strong generalization across datasets and domains. Extensive experiments on natural, aerial, and medical counting benchmarks validate its robustness and superior performance compared to SOTA methods.
@@ 2 Related Work
To eliminate the dependence on manually annotated exemplars in FSOC methods (Lu et al., 2018; Ranjan et al., 2021; Liu et al., 2022; DJukic et al., 2023; He et al., 2024), Xu et al. (Xu et al., 2023) introduce the ZOC task, which estimates object counts from a class name by employing a text-conditioned variational autoencoder (VAE) that synthesizes visual exemplars through patch–prototype pseudo-labels. However, the arbitrary selection of patches hinders precise object delineation, thereby degrading exemplar quality. Subsequent works (Jiang et al., 2023; Kang et al., 2024; Amini et al., 2023) leverage image–text alignment to model object–class correlations without relying on visual exemplars, enhancing scalability but often failing to preserve structural detail for irregularly shaped objects. Detection-driven approaches such as DAVE (Pelhan et al., 2024b) adopt a detect-and-verify pipeline but bias toward majority-class counts (Qian et al., 2025). Recent VLM-based frameworks like PseCo (Huang et al., 2024) employs a three-stage point, segment and count design integrating SAM (Kirillov et al., 2023) for segmentation and CLIP (Radford et al., 2021) for verification, while GeCo (Pelhan et al., 2024a) unifies detection, segmentation, and counting with a single-stage dense query encoder–decoder architecture. VA-Count (Zhu et al., 2024) instead leverages GroundingDINO for text-guided detection and performs a simple logits-based exemplar selection, yielding limited exemplar diversity and relying on DINO’s initial proposals (often multi-instance boxes). Our CountZES explicitly handles the challenge of limited or no single instance boxes coming from the detector and multi-stage exemplar selection under semantic, statistical, and feature-level priors results in diverse exemplars. Finally, T2ICount (Qian et al., 2025) exploits text-to-image diffusion priors for open-world counting but exhibits limited cross-domain generalization due to dataset-specific ZOC optimization (Figure. 1). In contrast, our CountZES being inference-only demonstrates superior generalization. Among the inference-only approaches TFOC (Shi et al., 2024), TFCAC (Lin et al., 2025), and OmniCount (Mondal et al., 2025) formulate counting as a segmentation problem and are constrained by the limited performance of SAM in handling small and congested objects (Ma et al., 2023).
Figure 2: (a) Detectors like GroundingDINO (Liu et al., 2024) often yield very few or even fail to provide single instance boxes. (b) Counting is sensitive to exemplar choice. (c) Introducing CountZES a multi-stage inference-only exemplar selection approach for ZOC. (Best viewed in zoom).
@@ 3 Method
Figure 3: Overview of CountZES. The pipeline operates in three stages. DAE refines GroundingDINO detections using CLIP similarity via SSES to obtain text-aligned single-instance regions. DGE then adopts a density-driven paradigm: it generates a density map conditioned on DAE exemplar, leverages P2P prompting to identify candidate exemplars, applies RoI-based single-instance filtering, and selects the most reliable exemplar through GGES. Finally, FCE focuses on visual coherence by projecting candidate exemplars onto SAM’s feature map and using FRES to select the representative exemplar. An exemplar from each stage form a diverse exemplar set for final object count.
Zero-shot exemplar discovery in dense scenes is inherently ambiguous due to semantic noise, multi-instance proposals, and appearance variability. To address these challenges, we introduce CountZES, an inference-only framework for ZOC (Figure. 3), which discovers a diverse exemplar set through three synergistic stages, namely DAE (Sec. 3.1), DGE (Sec. 3.2), and FCE (Sec. 3.3). DAE performs semantic grounding and single-instance isolation, DGE enforces statistical reliability via density-based self-supervision and pseudo-GT consistency, and FCE promotes visual representativeness through feature-level consensus.
@@ 3.1 Detection Anchored Exemplar (DAE)
Initial Box Selection: Each detection box $\mathbf{b}_{i}$ serves as a candidate region, within which $S$ provides fine-grained confidence cues about semantic alignment with $t$. However, not all detections correspond to clean, single-instance regions, as some may encompass multiple instances or include background areas. To quantify the internal consistency of each region, we require a measure that captures how concentrated or dispersed the similarity responses are within the box. We therefore compute the entropy of the CLIP similarity distribution inside $\mathbf{b}_{i}$, which reflects the spatial uniformity of similarity values (see Appendix. A.3 for details). Intuitively, boxes that include multiple objects or background exhibit high entropy due to fluctuating similarity responses across diverse semantic areas. In contrast, a box that tightly encloses a single, well-aligned instance yields low entropy, indicating coherent similarity responses in $S$.
@@
where $\alpha=0.5$ controls the trade-off between semantic confidence and spatial purity, the latter reflecting how consistent $S$ is within $\mathbf{b}_{i}$. Higher $J_{\text{box}}(\mathbf{b}_{i})$ indicates boxes that are both confident (high $c_{i}$) and internally consistent (low $\widehat{H}$), favoring single-instance regions over cluttered ones. The top-scoring box, $\mathbf{b}^{c}=\arg\max_{\mathbf{b}_{i}}J_{\text{box}}(\mathbf{b}_{i})$, is selected as the coarse region of interest. Why coarse? Although this entropy-weighted scoring favors regions with fewer overlapping instances or backgrounds, it does not guarantee that the selected box strictly contains a single instance. When all GroundingDINO detections correspond to grouped regions, even the top-scoring box may include multiple instances. Hence, this coarse selection is refined next by the SSES module to obtain an instance-specific region.
Sim-guided SAM-based Exemplar Selection (SSES): Given the coarse region $\mathbf{b}^{c}$, SSES refines it into a clean, instance-specific exemplar by leveraging the similarity map $S$ and SAM’s segmentation capability (see Figure. 9). The core intuition is that regions containing well-aligned object instances will exhibit distinct peaks in $S$, representing local maxima of semantic alignment with the target text. To identify high-similarity regions, we sample $k$ peaks from $S$ within $\mathbf{b}^{c}$ by applying percentile-based threshold rather than a fixed similarity cutoff. Percentiles are computed over the full similarity map $S$, normalizing thresholds to each image’s global similarity statistics. Starting from a high percentile (e.g., 90th), the threshold is gradually relaxed in fixed steps ($\Delta p=10$) until at most $k=16$ peaks are detected within $\mathbf{b}^{c}$ (see Appendix A.1). This adaptive relaxation dynamically adjusts the confidence level for peak selection based on the strength and density of responses in the coarse region. Formally, for a percentile $p$, the threshold is defined as $\tau(p)=\operatorname{Perc}_{p}(S)$, and we retain spatial locations inside $\Omega(\mathbf{b}^{c})=\{(x,y)\,|\,x_{0}\leq x\leq x_{1},\,y_{0}\leq y\leq y_{1}\}$ whose similarity values exceed this threshold:
@@
where $w_{\text{sim}}=w_{\text{ent}}=0.5$. The optimal mask is selected as $M_{\text{DAE}}=\arg\max_{M_{i}}J_{\text{mask}}(M_{i})$, and its corresponding bounding box ($\mathbf{b}_{\text{DAE}}$) defines the refined, instance-specific region. In essence, SSES acts as a fine-grained refinement stage, transforming the coarse detector-derived region $\mathbf{b}^{c}$ into a single-instance exemplar $\mathbf{b}_{\text{DAE}}$.